# Differentially Private Federated Learning: A Systematic Review

### v3

@@ 1. Introduction
In response, federated learning (FL) has emerged as a promising collaborative learning infrastructure. In FL systems, participants retain their data locally, eliminating the need to share sensitive raw data. Instead, they contribute to a collective learning process by training models locally and sharing only the model parameters with a central server. This parameter server aggregates the updates and redistributes the refined model to all participants. However, even though the raw data are kept locally, adversaries may still infer sensitive data from the shared model parameters, posing severe privacy concerns. For example, the contents of raw data can be inverted from the model parameters  (Zhu et al., 2019; Nasr et al., 2019; Wang et al., 2019a), or the membership information of the raw data can be inferred (Song et al., 2017; Melis et al., 2019). To enhance the privacy of FL systems, several methods have been proposed based on homomorphic encryption  (Damgård et al., 2012) or secure multiparty computation (MPC) (Mohassel and Zhang, 2017). But these techniques require significant computational overheads and do not protect the final output of the computation, leaving them vulnerable to privacy breaches (e.g., inference attacks). One of the state-of-the-art (SOTA) paradigms to mitigate privacy risks in FL is differential privacy (DP)  (Dwork et al., 2014). Many works (Feldman, 2020; Naseri et al., 2020; Jayaraman and Evans, 2019; Stock et al., 2022) have demonstrated that by adding proper DP noise during the local training phase or uploaded model parameters, can prevent unintentional leakage of private raw data (e.g., via the membership inference attacks (Song et al., 2017; Melis et al., 2019)).
Currently, federated learning with differential privacy has attracted significant interests (Naseri et al., 2020; Yang et al., 2023c; Xu et al., 2023a; Zhang et al., 2022). Apart from DP  [^1] DP also refer as centralized DP (CDP) to differ LDP in some paper. The “DP” mentioned in this paper can be seen as “CDP”., other DP models 222 We refer to DP, LDP, and the shuffle model collectively as DP models in our work. such as local differential privacy (LDP) and the shuffle model, which are derived from DP, are also widely used in FL (Zhao et al., 2020b; Yang et al., 2021; Liu et al., 2021a, 2023b). Many scholars are exploring the trade-offs between privacy protection, communication efficiency, and model performance. As the volume of publications on differentially private federated learning  [^3] In our paper, differentially private federated learning means federated learning with DP models, which can be DP, LDP or shuffle model. continues to grow, the task of summarizing and organizing this body of research has become both urgent and challenging. While there are existing surveys on privacy-preserving federated learning (Mothukuri et al., 2021; Yin et al., 2021), they mainly focus on secure multiparty computation or homomorphic encryption techniques. There are also survey studies for the differential privacy (Yang et al., 2023a; Wang et al., 2020d; Xiong et al., 2014, 2020), focusing on their applications in graph data (Mueller et al., 2022), medical data (Liu et al., 2023a), natural language processing (NLP) (Hu et al., 2023) or the industrial internet of things (IIOT) (Jiang et al., 2021), rather than the federated learning.
El et al. (El Ouadrhiri and Abdelhadi, 2022) presented a survey of differential privacy for deep and federated learning. Their work primarily focuses on centralized deep learning, with only a small portion discussing the LDP in FL. Recently, Zhang et al. (Zhang et al., 2023a) offered a review for differential privacy techniques in federated learning, but their discussion does not provide a comprehensive, technical, and systematic analysis of differential privacy within the federated learning context. Beyond that, our work differs from the above in several key aspects: 1) We propose a novel differentially private federated learning taxonomy, particularly categorize DP, LDP and shuffle model in FL, we focus on the guarantee and definition of differential privacy, which sets us apart from previous classifications based on the presence of a centralized trusted server (Zhang et al., 2023a; Naseri et al., 2020; Yang et al., 2023c; Wei et al., 2021). We contend that their classification approach lacks precision. Federated learning is inherently a composite framework; for instance, in a cross-silo setting [^4] According to the participating clients of FL, in cross-device setting, the number of clients is usually larger and consists of mobile devices. While clients are typically organizations or companies, and the number of clients is smaller in cross-silo setting (Kairouz et al., 2021b)., aside from a central server that aggregates model parameters, each client may also maintain a local server to manage data for training. Therefore, characterizing federated learning scenarios that lack a trusted central server for parameter aggregation solely as LDP is overly simplistic and inaccurate. 2) We elaborate on the relationship between DP, LDP, and the shuffle model, as well as their privacy guarantee in FL. 3) We not only explore DP models in horizontal FL but also delve into vertical FL and transfer FL.
@@
As show in Figure 1, our survey classifies differentially private federated learning based on two aspects: FL scenarios and DP models. In FL scenarios, we classify the works to Horizontal FL, Vertical FL and Transfer FL by the data partitioning setting. We have encompassed all scenarios of federated learning. Although the majority of works on differential privacy for FL has been conducted in the context of Horizontal FL, as more privacy issues are exposed in vertical FL and transfer FL (Pasquini et al., 2021; Jagielski et al., 2024), research on DP models in them is becoming increasingly important. In DP models, We discuss differentially private federated learning from a new perspective, specifically, We distinguish between (centralized) DP, LDP and shuffle model from the definition and privacy guarantee. While these models can generally be grouped under the broad spectrum of DP, there are strict distinctions when considering the specific definitions and privacy guarantee. The differentiation and relation between them will be described in detail in Section 2.2.4. Rigorous differentiation of DP models in differentially private federated learning, based on the definitions and privacy guarantee, is of significant importance. It not only clarifies the protection subjects of different DP models but also enhances the understanding of various implementation methods associated with these DP models. For DP and shuffle model, we further categorize into sample level and client level based on the definition of neighboring datasets – corresponding to the protected objects  [^5] The definition of neighboring datasets does not exist in LDP..
@@
  • Based on the above research discussion, we have proposed 5 promising directions for future research.
@@
The rest of the paper is organized as follows. Section 2 summaries the technical of federated learning and three DP models, and the fundamental mechanisms of differentially private FL. In Section 3, we discuss the various DP models in horizontal FL. Section 4 shows the differential privacy techniques in vertical FL and transfer FL. Section 5 further explores the applications of differentially private federated learning in the real world. Open challenges and future directions are introduced in Section 6, followed the conclusion in Section 7.
@@ 2. FL and DP Models
In this section, we will describe three FL scenarios, and explain the definition and relation of three DP models. Apart that, we will present basic properties and popular loss composition mechanisms in DP models. Lastly, we will introduce the fundamental used perturbation mechanisms in differentially private FL.
@@ 2.1. Federated Learning
Federated learning is a learning framework where multiple clients collaborate on a machine learning problem, coordinated by a central server  [^6] This is the core distinction with fully decentralized (peer-to-peer) learning (Kairouz et al., 2021b).. Instead of sharing or transferring raw data, clients keep their data locally and send model parameters for immediate aggregation to achieve the learning objective. Based on the characteristics of the data held by each participant, FL can be divided into the following categories (Li et al., 2021c).
@@ 2.1.1. Horizontal Federated Learning (HFL)
HFL refers to the FL scenarios where participants hold different samples while sharing the same feature space. The objective of HFL is to collaboratively train a global model $\mathbf{w}^{\star}$ with all clients, aiming to minimize the global objective function $F\left(\cdot\right)$, essentially seeking to find:
@@
Here, $t\in[T]$ donates the communication rounds, $k\in[K]$ donates the clients, $L_{k}(\cdot)$ donates the loss function of client $k$, $\mathbf{w}^{k}_{t}$ signifies the model uploaded by client $k$ at round $t$, while $\mathbf{w}_{t}$ denotes the server aggregated model. The weights are defined as $p_{k}=\frac{|D_{k}|}{|D|}$, $\sum_{k=1}^{K}p_{k}=1$, $D_{k}$ represents the local dataset of client $k$, and $D$ is the union of $D_{k}$.

### v4

@@ 1. Introduction
In response, federated learning (FL) has emerged as a promising collaborative learning infrastructure. In FL systems, participants retain their data locally, eliminating the need to share sensitive raw data. Instead, they contribute to a collective learning process by training models locally and sharing only the model parameters with a central server. The central server aggregates the updates and redistributes the refined model to all participants. However, even though the raw data are kept locally, adversaries may still infer sensitive data from the shared model parameters, posing severe privacy concerns. For example, the contents of raw data can be inverted from the model parameters  (Zhu et al., 2019; Nasr et al., 2019; Wang et al., 2019a), or the membership information of the raw data can be inferred (Song et al., 2017; Melis et al., 2019). To enhance the privacy of FL systems, several methods have been proposed based on homomorphic encryption  (Damgård et al., 2012) or secure multiparty computation (MPC) (Mohassel and Zhang, 2017). But these techniques require significant computational overheads and do not protect the final output of the computation, leaving them vulnerable to privacy breaches (e.g., inference attacks). One of the state-of-the-art (SOTA) paradigms to mitigate privacy risks in FL is differential privacy (DP)  (Dwork et al., 2014). Many works (Feldman, 2020; Naseri et al., 2020; Jayaraman and Evans, 2019; Stock et al., 2022) have demonstrated that by adding proper DP noise during the local training phase or uploaded model parameters, can prevent unintentional leakage of private raw data (e.g., via the membership inference attacks (Song et al., 2017; Melis et al., 2019)).
Currently, federated learning with differential privacy has attracted significant interests (Naseri et al., 2020; Yang et al., 2023b; Xu et al., 2023a; Zhang et al., 2022). As the volume of publications on differentially private federated learning continues to grow, the task of summarizing and organizing this body of research has become both urgent and challenging. Despite the existence of surveys on differentially private federated learning (El Ouadrhiri and Abdelhadi, 2022; Zhang et al., 2023a; Farooqi et al., 2024; Ren et al., 2024), they do not provide a comprehensive, technical, and systematic analysis of differential privacy within the federated learning context. Table 1 demonstrates the distinctions between our work and existing surveys. Specifically, our work differs from their in several key aspects: 1) We not only explore DP models in horizontal FL but also delve into vertical FL and transfer FL; 2) We propose a novel differentially private federated learning taxonomy based on the guarantees and definitions of DP, specifically categorizing DP [^1] We note that previous literature often employs CDP (Centralized DP) in FL to distinguish it from LDP. The definition of CDP in those works is consistent with the definition of DP used in our work. In this work, we deliberately retain the original notation DP for differential privacy in the centralized setting, instead of introducing the additional symbol CDP., local DP (LDP), and the shuffle model in federated learning. Furthermore, we dissect the interconnections among these DP models and their respective privacy guarantees within the federated learning framework; 3) We elaborate on the relationship between DP, LDP, and the shuffle model in FL, and investigate different neighborhood levels for each DP models; 4) We provide a elaboration on different differential privacy composition mechanisms and their resulting effects; 5) We demonstrate the applications of differentially private federated learning in real-world scenarios.
Table 1. The comparison between related works and ours.
Releated works Federated scenario DP model Neighborhood level Composition mechanism Application
Horizontal Vertical Transfer CDP (DP) LDP Shuffle
Farooqi et al. (Farooqi et al., 2024)
Ren et al. (Ren et al., 2024)
El et al. (El Ouadrhiri and Abdelhadi, 2022)
Zhang et al. (Zhang et al., 2023a)
Ours
@@
As show in Figure 1, our survey classifies differentially private federated learning based on two aspects: FL scenarios and DP models. In FL scenarios, we classify the works to Horizontal FL, Vertical FL and Transfer FL by the data partitioning setting. We have encompassed all scenarios of federated learning. Although the majority of works on differential privacy for FL has been conducted in the context of Horizontal FL, as more privacy issues are exposed in vertical FL and transfer FL (Pasquini et al., 2021; Jagielski et al., 2024), research on DP models in them is becoming increasingly important. In DP models, We discuss differentially private federated learning from a new perspective, specifically, We distinguish between DP, LDP and shuffle model from the definition and privacy guarantee, which sets us apart from previous classifications based on the presence of a centralized trusted server (Zhang et al., 2023a; Naseri et al., 2020; Yang et al., 2023b; Wei et al., 2021). We contend that their classification approach lacks precision. Federated learning is inherently a composite framework; for instance, in a cross-silo setting [^2] According to the participating clients of FL, in cross-device setting, the number of clients is usually larger and consists of mobile devices. While clients are typically organizations or companies, and the number of clients is smaller in cross-silo setting (Kairouz et al., 2021b)., aside from a central server that aggregates model parameters, each client may also maintain a local server to manage data for training. Therefore, characterizing federated learning scenarios that lack a trusted central server for parameter aggregation solely as LDP is overly simplistic and inaccurate. While these models can generally be grouped under the broad spectrum of DP, there are strict distinctions when considering the specific definitions and privacy guarantee. The differentiation and relation between them will be described in detail in Section 2.2.4. Rigorous differentiation of DP models in differentially private federated learning, based on the definitions and privacy guarantee, is of significant importance. It not only clarifies the protection subjects of different DP models but also enhances the understanding of various implementation methods associated with these DP models. For DP and shuffle model, we further categorize into sample level and client level based on the definition of neighboring datasets – corresponding to the protected objects  [^3] The definition of neighboring datasets does not exist in LDP..
@@
  • Based on the above research discussion, we have proposed 6 promising directions for future research.
@@
The rest of the paper is organized as follows. As shown in Table 2, we will first introduce the research status of three DP models in horizontal federated learning, followed by the three DP models in vertical federated learning, and finally the three DP models in transfer federated learning. Specially, Section 2 presents three FL settings and three DP models. In Section 3, we discuss the various DP models in horizontal FL. Section 4 shows the differential privacy techniques in vertical FL and transfer FL. Section 5 further explores the applications of differentially private federated learning in the real world. Open challenges and future directions are introduced in Section 6, followed the conclusion in Section 7.
Table 2. An overview study of Differentially Private FL.
Federated
Scenario
Publications Year
DP
Model
Neighborhood
Level
Perturbation
Mechanism
$\textbf{CM}^{1}$
Downsteam
Tasks
$\begin{tabular}[c]{@{}c@{}}Model\\ Archiecture\end{tabular}^{2}$
Clients
Number
$\epsilon$ $\delta$
Chen et al.(Chen et al., 2024) 2024 Gaussian tCDP Classification LR, Shallow CNN 100 0.3 $10^{-2}$
malekmohammadi et al. (Malekmohammadi et al., 2024) 2024 Gaussian AC Classification CNN [20,60] [0.5,5] $10^{-4}$
Liu et al.(Liu et al., 2024b) 2024 Gaussian RDP Classification CNN 10 [0.1,10] $10^{-3}$
Ling et al. (Ling et al., 2024) 2024 Gaussian RDP Classification Shallow CNN 10 [1.5,5.5] $10^{-5}$
Xiang et al. (Xiang et al., 2023) 2023 Gaussian MA Classification Shallow CNN,LSTM [10,20] [0.12,2] $10^{-2},10^{-5}]$
Ruan et al. (Ruan et al., 2023) 2023 Gaussian RDP Classification Shallow CNN, LSTM [3,10] [0.25,2] $10^{-4},10^{-5}]$
Noble et al. (Noble et al., 2022) 2022 Gaussian RDP Classification Shallow CNN 10 [3,13] $10^{-6}$
Fu et al. (Fu et al., 2022a) 2022 Gaussian RDP Classification Shallow CNN 10 [2,6] $10^{-5}$
Li et al. (Li et al., 2022c) 2022 Gaussian MA Classification LR, Shallow CNN 10 [1,16] $10^{-3}$
Ryu et al. (Ryu and Kim, 2022) 2022 Gaussian AC Classification LR [10,195] [0.05,5] $10^{-6}$
Wei et al. (Wei et al., 2021) 2021 Gaussian MA Classification Shallow CNN 50 [4,20] $10^{-3}$
Liu et al. (Liu et al., 2021b) 2021 Gaussian GDP Classification Shallow CNN 100 [10,100] $10^{-3}$
Zheng et al. (Zheng et al., 2021) 2021 Gaussian GDP Classification Shallow CNN 100 [10,100] $10^{-3}$
Huang et al. (Huang et al., 2020) 2020 Gaussian, Laplace AC Classification Shallow CNN 10,100,1000 [0.2,8] $10^{-2},10^{-5}]$
Wei et al. (Wei et al., 2020) 2020 Gaussian MA Classification Shallow CNN,LSTM [10,20] [0.12,2] $10^{-2},10^{-5}]$
Huang et al. (Huang et al., 2019) 2019 SL Gaussian AC Regression LR - [0.01,0.2] $10^{-3},10^{-6}]$
Yang et al. (Yang et al., 2023a) 2023 Gaussian RDP Classification Shallow CNN 50 [2,16] $10^{-3}$
Xu et al. (Xu et al., 2023a) 2023 Gaussian RDP Classification ResNet-50 [1262,9896000] [10,20] $10^{-7}$
Shi et al. (Shi et al., 2023) 2023 Gaussian RDP Classification ResNet-18 500 [4,10] $\frac{1}{500}$
Zhang et al. (Zhang et al., 2022) 2022 Gaussian MA Classification Shallow CNN, ResNet-18 1920 [1.5,5] $10^{-5}$
Cheng et al. (Cheng et al., 2022) 2022 Gaussian MA Classification Shallow CNN, ResNet-18 3400 [2,8] $\frac{1}{3400}$
Bietti et al. (Bietti et al., 2022) 2022 Gaussian MA Classification Shallow CNN 1000 [0.1,1000] $10^{-4}$
Andrew et al. (Andrew et al., 2021) 2021 Gaussian RDP Classification Shallow CNN [500,342000] [0.035,5] $\frac{1}{500},\frac{1}{342000}]$
Mamahan et al. (McMahan et al., 2017b) 2017 Gaussian MA Classification LSTM [100,763430] [2.0,4.6] $10^{-9}$
Geyer et al. (Geyer et al., 2017) 2017 CL Gaussian MA Classification Shallow CNN 100, 1000, 10000 8 $10^{-3},10^{-6}]$
Chen et al. (Chen et al., 2022a) 2022 Discrete Gaussian RDP Classification Shallow CNN [100,1000] [0,10] $10^{-2}$
Chen et al. (Chen et al., 2022b) 2022 Poisson Binomial RDP Classification LR 1000 [0.5,6] $10^{-5}$
Wang et al. (Wang et al., 2020a) 2020 Discrete Gaussian RDP Classification Shallow CNN 100K [2,4] $10^{-5}$
Stevens et al. (Stevens et al., 2022) 2022 LWE RDP Classification Shallow CNN [500,1000] [2,8] $10^{-5}$
Kairouz et al. (Kairouz et al., 2021a) 2021 Discrete Gaussian zCDP Classification Shallow CNN 3400 [3,10] $\frac{1}{3400}$
Agarwal et al. (Agarwal et al., 2021) 2021 Skellam RDP Classification Shallow CNN 1000k [5,20] $10^{-6}$
Kerkouche et al. (Kerkouche et al., 2021) 2021 Gaussian MA Classification Shallow CNN [5011,6000] [0.5,1] $10^{-5}$
Agarwal et al. (Agarwal et al., 2018) 2018 CL with SA Binomial AC Classification LR 25M [2,4] $10^{-9}$
Naseri et al. (Naseri et al., 2020) 2020 SL, CL Gaussian RDP Classification Shallow CNN,LSTM [100,660120] [1.2,10.7] $10^{-5}$
Yang et al. (Yang et al., 2023b) 2023 DP SL, CL, CL with SA Gaussian, Skellam RDP Classification Shallow CNN [40,500] [2,8] $10^{-3}$
Triastcyn et al. (Triastcyn and Faltings, 2019) 2019 Bayesian DP SL, CL Gaussian RDP Classification ResNet-50 [100,10000] [0.2,4] $10^{-3},10^{-6}]$
Zhang et al. (Zhang et al., 2024) 2024 Gaussian zCDP Classification LR 20 1 $10^{-4}$
Varun et al. (Varun et al., 2024) 2024 SRR BC Classification Shallow CNN 100 [1,10] 0
Zhang et al. (Zhang et al., 2023b) 2023 Gaussian AC Classification LR, Shallow CNN 100 [3,30] -
Wang et al. (Wang et al., 2023a) 2023 EM, DMP-UE BC Classification Shallow CNN [10,50] [0.1,1] 0
Jiang et al. (Jiang et al., 2022) 2023 EM BC Classification Shallow CNN [100,750] [0.5,12] 0
Li et al. (Li et al., 2022b) 2022 Laplace BC Classification Shallow CNN 100 78.5 0
Lian et al. (Lian et al., 2022) 2022 Laplace BC Classification Shallow CNN 5 [3,6] 0
Mahawaga et al. (Mahawaga Arachchige et al., 2022) 2022 RAPPOR BC Classification Shallow CNN [2,100] [0.5,10] 0
Wang et al. (Wang et al., 2022b) 2022 RAPPOR BC Classification LR [500,1800] [0.1,10] 0
Zhao et al. (Zhao et al., 2022) 2022 Adaptive-Harmony BC Classification Shallow CNN 200 [1,10] 0
Sun et al. (Sun et al., 2021) 2021 Adaptive-Duchi BC Classification Shallow CNN [100,500] [1,5] 0
Yang et al. (Yang et al., 2021) 2021 Laplace BC Classification Shallow CNN [200,1000] [1,10] 0
Wang et al. (Wang et al., 2020c) 2020 RRP AC Topic Modeling LDA 150 [5,8] [0.05,0.5]
Zhao et al. (Zhao et al., 2020b) 2020 Three output, PM-SUB BC Classification LR, SVM 4M [0.5,4] 0
Liu et al. (Liu et al., 2020a) 2020 RR, PM BC Classification LR, SVM 4W-10W [0.5,16] 0
Wang et al. (Wang et al., 2019b) 2019 LDP - PM BC Classification LR, SVM 4M [0.5,4] 0
Truex et al. (Truex et al., 2020) 2020 Condensed LDP - EM BC Classification Shallow CNN 50 1 0
Liu et al. (Liu et al., 2023) 2023 Clipped-Laplace,Shuffle AC Classification LR 10000 25.6 $10^{-8}$
Liew et al. (Liew et al., 2022) 2022 Harmony,Shuffle RDP Classification Shallow CNN [50000,60000] [2.8] -
Liu et al. (Liu et al., 2021a) 2021 CL Laplace,Shuffle BC, AC Classification LR 1000 4.696 $5*10^{-6}$
Chen et al. (Chen et al., 2023a) 2023 Duchi,Shuffle GDP Classification Shallow CNN 100 [0.5, 100] $10^{-5}$
Horizontal Girgis et al. (Girgis et al., 2021b) 2021 Shuffle DP SL Laplace,Shuffle AC Classification Shallow CNN 60000 [1,10] $10^{-5}$
Takahashi et al. (Takahashi et al., 2023) 2023 KRR BC Classification GBDT 3 [0.1,2.0] -
Yang et al. (Yang et al., 2022) 2022 Label DP Laplace, KRR BC Classification Shallow CNN 2 1 0
Oh et al. (Oh et al., 2022) 2022 SL Gaussian RDP Classification VGG-16 10 [1,40] -
Chen et al. (Chen et al., 2020) 2020 Gaussian GDP Classification Shallow CNN [3,8] - -
Wang et al. (Wang et al., 2020b) 2020 Gaussian AC Classification Shallow CNN 2 [0.001, 10] $10^{-2}$
Wu et al. (Wu et al., 2020) 2020 DP CL Laplace BC Classification GBDT [2,10] - -
Mao et al. (Mao et al., 2022) 2022 Laplace, RR BC Classification Shallow CNN 5 [0.1,4.0] 0
Tian et al. (tian2020federboost) 2020 LDP - $R$ BC Classification GBDT 3 4 0
Vertical Li et al. (Li et al., 2022a) 2022 Condensed LDP - Discrete Laplace BC Classification GBDT 2 [0.64,2.56] 0
Wan et al. (Wan et al., 2023) 2023 Gaussian AC Recommendection DeepFM 2 [0.05, 10] -
Hoech et al. (Hoech et al., 2022) 2022 Gaussian AC Classification Resnet-18 20 [0.1,0.5] -
Tian et al. (Tian et al., 2022) 2022 Gaussian GDP Text Generation GPT-2 2000 [3,5] $10^{-6}$
Sun et al. (Sun and Lyu, 2020) 2020 Random Sampling AC Classification Shallow CNN 6 [0.003,0.65] [0.006,0.65]
Papernot er al. (Papernot et al., 2018) 2018 Gaussian RDP Classification Resnet-18 2 [0.59,8.03] $10^{-8}$
Papernot er al. (Papernot et al., 2017) 2017 SL Laplace MA Classification Shallow CNN 2 [2.04,8.19] $10^{-5},10^{-6}]$
Dodwadmath et al. (Dodwadmath and Stich, 2022) 2022 Laplace MA Classification Shallow CNN 10 [11.75,20] $10^{-5}$
Pan et al. (Pan et al., 2021a) 2021 DP CL Gaussian RDP Classification Resnet-18 100 [0.95,9.03] -
Transfer Qi et al. (Qi et al., 2023) 2023 LDP - KRR BC Classification Shallow CNN [2,5] [2,7] 0
  • 1
    1. CM=Composition Mechanism, BC=Basic Sequential Composition Theory, AC=Advanced Sequential Composition Theory.
  • 2
@@ 2. FL and DP Models
In this section, we will describe three FL scenarios, and explain the definition and relation of three DP models. Additionally, we provide a discussion on popular loss composition mechanisms in DP models and fundamental used perturbation mechanisms for differentially private FL in Section A and Section B, respectively.
@@ 2.1. Federated Learning
Federated learning (FL) is a distributed machine learning paradigm where multiple clients collaboratively train a global model under the coordination of a central server, without sharing their raw data. Based on how data is distributed among participants (Li et al., 2021c), FL can be categorized as follows.
@@ 2.1.1. Horizontal Federated Learning (HFL)
HFL applies to scenarios where participants hold different data samples but share the same feature space. Its objective is to collaboratively train a global model $\mathbf{w}^{\star}$ by aggregating model updates from all clients to minimize a weighted global loss function $F(\mathbf{w})$:
@@
Here, $p_{k}$ is the aggregate weight of client $k$, $L_{k}(\cdot)$ is its loss function, and $\mathbf{w}_{t}^{k}$ is the model uploaded by client $k$ at communication round $t$.